OpenStudy (anonymous):

use the laws of exponents to simplify use the laws of exponents to simplify @Mathematics

OpenStudy (anonymous):

\[2_{9}^{6} \over 2^{_{7}^{6}}\]

OpenStudy (anonymous):

that is 2^9/6 over 2^7/6

OpenStudy (saifoo.khan):

2^16/6 = 2^8/3

OpenStudy (saifoo.khan):

have more?

OpenStudy (anonymous):

yes if you want to do some more for me

OpenStudy (saifoo.khan):

sure.

OpenStudy (anonymous):

simplify by factoring \[\sqrt{700x ^{4}}\]

OpenStudy (saifoo.khan):

\[10x^2 \sqrt7\]

OpenStudy (anonymous):

add simplify by collecting like radical terms \[17\sqrt{48}+3\sqrt{108}\]

OpenStudy (saifoo.khan):

\[86 \sqrt3\]

OpenStudy (anonymous):

multiply \[\sqrt[3]{2}\times \sqrt[3]{4}\] that is 3squart 2 times 3squart 4

OpenStudy (anonymous):

saifoo shouldn't the first one be \[2^{\frac{1}{3}}\]?

OpenStudy (saifoo.khan):

2

OpenStudy (saifoo.khan):

OMG wait. hr had a different question before. he changed it when.. idk.

OpenStudy (anonymous):

use rational exponents to simplify write the answer in radicnal notation \[ \sqrt[12]{a ^{10}}\]

OpenStudy (saifoo.khan):

\[\sqrt[6]{a^5}\]

OpenStudy (anonymous):

so my last question saifoo you wrote 2 is that the answer to my last questions?

OpenStudy (saifoo.khan):

\[\sqrt[3]{2}\times \sqrt[3]{4} = 2\]

OpenStudy (anonymous):

=what is the number before the squart root of a^5

OpenStudy (saifoo.khan):

six

OpenStudy (anonymous):

the formula \[r=2\sqrt{5L}\] can be used to approximate the speed r in miles per hour of a car that has left skid marks of length L, in feet how far will a car skid at 50 mph

OpenStudy (saifoo.khan):

125

OpenStudy (anonymous):

multiply \[\sqrt[4]{16} \times \sqrt[4]{32}\]

OpenStudy (saifoo.khan):

\[\Huge4 \sqrt[4]{2}\]

OpenStudy (anonymous):

are you done or can you help with more

OpenStudy (saifoo.khan):

let me guess, u have to submit that before 12?

OpenStudy (anonymous):

haha ya

OpenStudy (saifoo.khan):

UoP ? mathlab?

OpenStudy (anonymous):

i have this and 4 others and a final

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

u go to the same one

OpenStudy (saifoo.khan):

just 4 out of 66? :O

OpenStudy (anonymous):

just four out of 66 what?

OpenStudy (saifoo.khan):

No, i help manyyy UoP guys with their hw. i have done these problems millions time.

OpenStudy (saifoo.khan):

times*

OpenStudy (anonymous):

o ok

OpenStudy (saifoo.khan):

final has 66 questions

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i havent even start that yet this is for my quiz

OpenStudy (saifoo.khan):

ohh.

OpenStudy (anonymous):

add simplify by collecting like radical terms if possible assume that all expressions under radicals represent non-negative numbers \[3\sqrt{y}+6\sqrt{y}\]

OpenStudy (anonymous):

you can stay on and help with the final if you like lol

OpenStudy (saifoo.khan):

\[9 \sqrt y\] lol. 66 questions. are u srs?

OpenStudy (anonymous):

simplify \[\sqrt[3]{-729x ^{3}}\]

OpenStudy (anonymous):

you sound smart if you get al these right i will know when i submit it

OpenStudy (saifoo.khan):

-9x

OpenStudy (saifoo.khan):

lol, u want "all" correct.. i dont take gurantee for one or two.. rest all are correct.

OpenStudy (anonymous):

subtract simplify by collectin glike radical terms if possible assume all expressions under radicals represent non-negative numbers \[\sqrt[3]{1250x}-\sqrt[3]{10x ^{4}}\]

OpenStudy (anonymous):

lol as long as i get a passing grade and i dont fault the quizzes ol i dont need all them correct

OpenStudy (saifoo.khan):

lol, u will get 80% + . that's for sure.

OpenStudy (anonymous):

im here interent messed up

OpenStudy (anonymous):

did you see i posted another one would you be able to work that one?

OpenStudy (saifoo.khan):

this site is lagging. :O

OpenStudy (anonymous):

what does it mean to refer to a 20-in tv or 25-in tv set? such units refer to the diagonal of the screen a 15-in TV set also has a width of 12 inches what is it height?

OpenStudy (saifoo.khan):

9

OpenStudy (anonymous):

if you still would like to help me

OpenStudy (saifoo.khan):

umm, Alright =D

OpenStudy (saifoo.khan):

this is quiz right?

OpenStudy (anonymous):

this is the squiz not my final i have posted two questions could you help me with

OpenStudy (saifoo.khan):

where are those two?

OpenStudy (anonymous):

scroll up can you not see them'/

OpenStudy (anonymous):

about the tv

OpenStudy (anonymous):

what does it mean to refer to a 20-in tv or 25-in tv set? such units refer to the diagonal of the screen a 15-in TV set also has a width of 12 inches what is it height?

OpenStudy (saifoo.khan):

what does it mean to refer to a 20-in tv or 25-in tv set? such units refer to the diagonal of the screen a 15-in TV set also has a width of 12 inches what is it height? 9

OpenStudy (anonymous):

subtract simplify by collect like radical terms\[\sqrt[3]{1250x}-\sqrt[3]{10x ^{4}}\]

OpenStudy (anonymous):

that is 3 in front of the 1250 x and 3 squar 10x^4

OpenStudy (saifoo.khan):

\[\sqrt[3]{10x}(5-x)\]

OpenStudy (anonymous):

simplify by taking roots of numberators and denomiator \[\sqrt{25 \over 121}\]

OpenStudy (saifoo.khan):

121 or 125?

OpenStudy (anonymous):

i still have 5 on the quiz if you wouldnt mind helpint with that

OpenStudy (anonymous):

121

OpenStudy (anonymous):

and myabe some on my final if you hvae time it not due till sunday maybe i will catch you on again tomorrow or sunday you can help me more?

OpenStudy (saifoo.khan):

that cannot be simplified.. well sure.. u may email me.. or facebook me.. we can talk about the final. ^_^

OpenStudy (anonymous):

yes i hate math i used to be good at it and i got so behind i have to have all this turned in by midnight and it sucks

OpenStudy (anonymous):

solve \[\sqrt{5y+6}=4\]

OpenStudy (saifoo.khan):

2

OpenStudy (anonymous):

divide then simplify by taking roots if possible \[\sqrt{9xy ^{3}} \over \sqrt{3x}\]

OpenStudy (anonymous):

that is 9xy ^3 and then 3x on bottom

OpenStudy (saifoo.khan):

\[y \sqrt3\]

OpenStudy (anonymous):

find the following \[\sqrt{(a+7)^{2}}\]

OpenStudy (anonymous):

that is ^2 at the end

OpenStudy (saifoo.khan):

|a+7|

OpenStudy (anonymous):

solv\[\sqrt[3]{5x+9}+9=6\]e

OpenStudy (saifoo.khan):

no solution

OpenStudy (anonymous):

simplify \[4^{3 \over 2}\]

OpenStudy (anonymous):

that is 4 then ^ 3 over 2

OpenStudy (saifoo.khan):

8

OpenStudy (anonymous):

i got 72. 92 thanks

OpenStudy (saifoo.khan):

dang! that's a C. :O

OpenStudy (saifoo.khan):

must be some format mistakes!

OpenStudy (anonymous):

its ok better than what i would have gotten

OpenStudy (anonymous):

now its my final i have to wrry about

OpenStudy (saifoo.khan):

do u have an email id?

OpenStudy (anonymous):

my email personal email is brittanigraham88@yahoo.com

OpenStudy (saifoo.khan):

let's talk on email.

OpenStudy (anonymous):

ok that is fine u wan tto help now with final i have a facebook give me yoru email for facebook that is easier

OpenStudy (saifoo.khan):

einstein.saifoo@gmail.com

OpenStudy (anonymous):

added